A block of metal weighing 2kg is resting on a frictionless plane. It is struck by a jet releasing water at a rate of 1kg/s a speed of 5m/s. The initial acceleration of the block will

Grade:12

1 Answers

Bhartu
6 years ago
F=vdm/dt..........(1)F=ma.....(2)HenceFrom 1 and 2Vdm/dt=ma 5*1=2*a5/2=2.5m/s2The acceleration will be 2.5 metre per second square.!